OpenStudy (anonymous):

What is 3/8 * 5/6? Will give medal!! A)9/20 B)4/7 C)5/16 D)1/6

OpenStudy (anonymous):

first response gets medal

OpenStudy (eli_moses):

C

OpenStudy (anonymous):

aye.

OpenStudy (anonymous):

it is C!!

OpenStudy (anonymous):

\[\dfrac{3}{8} * \dfrac{5}{6} = \dfrac{15}{48} -> \dfrac{5}{}\]

OpenStudy (anonymous):

ok let me see what i get and whoever got the same answer gets the medal